OpenStudy (anonymous):

Please help. How do I start this? Radium-221 has a half-life of 30 sec. How long will it take for 96% of a sample to decay?

OpenStudy (anonymous):

First you have to find how many half-lifes it will take for 96% of a sample to decay. 96% to decay meaning you are left with 4% of the sample, equivalent to 4/100 = 1/25. (1/2) ^ n = 1/25 Where n = number of half-lives. To explain that above equation. note that when n=1, after one half-life, you decay half of it leaving 1/2. When n=2, after 2 half-life, (1/2)^2 = 1/4. Thus if you notice the above equation is actually stating the amount of sample i would get after n amount of half-lifes. Solve for n in that equation. (1/2) ^ n = 1/25 log 2 ^ n = log 25 n = log 25/ log 2 = 4.644 ( 4sf). Time taken = number of half-live * duration of a half-life = 4.644 * 30 = 139.32 = 139 sec ( 3 sf) approx.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

Np.